OpenStudy (anonymous):

[9.01] When the quadratic equation y = (x − 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"? 2 3 −3 −20

OpenStudy (cwrw238):

expand (x − 4)(2x + 5) to get it in the form ax^2 + bx + c = 0 first multiply (2x + 5) by x then by -4 and simplify

OpenStudy (anonymous):

thank you but i dont understand how u expand it

OpenStudy (cwrw238):

* in the form y = ax^2 + bx + c

OpenStudy (anonymous):

oh i think i undersatnd now the answer i got was -3 is that right?

OpenStudy (cwrw238):

first multiply 2x + 5 by x: = x*2x + x*5 = 2x^2 + 5x then by -4:- = -4*2x +(-4*5) can you finish this part?

OpenStudy (cwrw238):

first multiply 2x + 5 by x: = x*2x + x*5 = 2x^2 + 5x then by -4:- = -4*2x +(-4*5) can you finish this part?

OpenStudy (anonymous):

(8x+-20)?

OpenStudy (anonymous):

nevermind i did it wrong .. lemme start over

OpenStudy (anonymous):

2x^2+5x then what happens after?

OpenStudy (anonymous):

how do you get the: then by -4:- = -4*2x +(-4*5) can you finish this part? i dont understand that

OpenStudy (cwrw238):

you had the signs wrong in second part - time + - so it should be -8x - 20

OpenStudy (cwrw238):

you had the signs wrong in second part - time + - so it should be -8x - 20

OpenStudy (anonymous):

oh i see then what comes after?

OpenStudy (cwrw238):

ok so now you have 2x^2 + 5x - 8x - 20 now you can add the 2 middle terms terms + 5x - 8x = -3x so it simplifies to y = 2x^2 - 3x - 20 so now you have the coefficients a b and c

OpenStudy (anonymous):

ohhh! i get it now and "b" is -3! thank u so much :D that helped alot

OpenStudy (cwrw238):

welcome